(a) What is the approximate force of gravity on a 70-kg person due to the Andromeda galaxy, assuming its total mass is 1013 that of our Sun and acts like a single mass 2 Mly away?
1
Expert's answer
2020-04-06T09:20:23-0400

1 Mly ="9.45\times10^{21}m"

2 Mly ="18.9\times10^{21}m"

mass of sun ="1.989 \u00d7 10^{30} kg"

mass of  Andromeda="1013\times1.989\times10^{30}=2\times10^{31}kg"

Force of gravity is given by "G\frac{Mm}{r^2}"

Force = "6.67259 \times 10^{-11}\times\frac{2\times10^{31}\times70}{(1.89\times10^{30})^2}=2.61\times10^{-40}N"